OpenStudy (anonymous):

If the ratio of N to K is 7 to 8 and the ratio of R to K is 6 to 11, what is the ratio of R to N? (a) 48/77 (b) 6/7 (c) 77/88 (d) 48/88 (e) 77/48

OpenStudy (lgbasallote):

is this figure-related?

OpenStudy (lgbasallote):

nevermind

OpenStudy (lgbasallote):

\[\frac NK = \frac 78\] and \[\frac RK = \frac{6}{11}\] right?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

so \[\huge \frac{\frac{R}{K}}{\frac NK} = \frac RN\] agree?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so can you solve the ratio of R to N now?

OpenStudy (anonymous):

is it (a) 48/77?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome :D